# The quadratic linking degree

### v5

@@ 1. Introduction
We begin by defining counterparts, over a perfect field $F$, to oriented links with two components (i.e. couples of disjoint oriented knots). Specifically, we replace the circle $\mathbb{S}^{1}$ with $\mathbb{A}^{2}_{F}\setminus\{0\}$ and the $3$-sphere $\mathbb{S}^{3}$ with $\mathbb{A}^{4}_{F}\setminus\{0\}$. We then define a counterpart to the linking number, which in knot theory is an invariant of oriented links with two components which corresponds to the number of times one of the oriented components turns around the other oriented component, and a counterpart to the linking couple, which is a couple of integers whose absolute values coincide with the absolute value of the linking number. We call these counterparts the ambient quadratic linking degree and the quadratic linking degree (couple) respectively. To define these, we use Chow-Witt groups, and more generally Rost-Schmid groups, instead of the singular cohomology groups used in classical knot theory. The Rost-Schmid groups are the cohomology groups of the Rost-Schmid complex (see Appendix B) which is built from Milnor-Witt $K$-theory groups; this is where motivic homotopy theory comes into play, as Milnor-Witt $K$-theory groups are in fact groups of morphisms of motivic spheres in the stable motivic homotopy category (see [Mor12, Corollary 1.25]).
@@
In Section 2, we give the definitions of the quadratic linking class, the quadratic linking degree and the ambient quadratic linking degree, then we determine how they depend on choices of orientations and of parametrisations of $\mathbb{A}^{2}_{F}\setminus\{0\}\to\mathbb{A}^{4}_{F}\setminus\{0\}$ (see Lemma 2.16) and deduce invariants of the (ambient) quadratic linking degree (see Corollaries 2.17 and 2.18 and Theorem 2.21). For instance, in the case $F=\mathbb{R}$, the absolute values of the components of the quadratic linking degree and of the ambient quadratic linking degree (which are in $\operatorname{W}(\mathbb{R})\simeq\mathbb{Z}$) are invariant under changes of orientations and of parametrisations of $\mathbb{A}^{2}_{F}\setminus\{0\}\to\mathbb{A}^{4}_{F}\setminus\{0\}$. This is similar to the fact that the absolute value of the linking number does not depend on choices of orientations. In the general case, the ranks modulo $2$ of the components of the quadratic linking degree and of the ambient quadratic linking degree are invariants, and more importantly we have the following lemma-definition and theorem:
@@
In Section 3 we give a method to explicitly compute the quadratic linking class (see Theorem 3.1), the quadratic linking degree (see Theorem 3.2) and the ambient quadratic linking degree (see Theorem 3.3) when the link $Z_{1}\sqcup Z_{2}\subset\mathbb{A}^{4}_{F}\setminus\{0\}$ is such that for each $i\in\{1,2\}$ the closure $\overline{Z_{i}}\subset\mathbb{A}^{4}_{F}$ of $Z_{i}$ is given by two irreducible equations $\{f_{i}=0,g_{i}=0\}$ such that $\{g_{1}=0,g_{2}=0\}$ is of codimension $2$ in $(\mathbb{A}^{4}_{F}\setminus\{0\})\setminus(Z_{1}\sqcup Z_{2})$ and for each generic point $p$ of an irreducible component of $\{g_{1}=0,g_{2}=0\}$, $f_{1}$ and $f_{2}$ are units in the residue field $\kappa(p)$.
@@ 2.1. Conventions and notations
where $\partial$ is the boundary map (see Definition B.12 and Theorem B.13) and the map in the middle is induced by the isomorphism $\det(\mathcal{N}_{\{0\}/\mathbb{A}^{2}_{F}})\to\mathcal{O}_{\{0\}}\otimes\mathcal{O}_{\{0\}}$ which sends $\overline{u}^{*}\wedge\overline{v}^{*}$ to $1\otimes 1$ (see Definition B.18). We also identify $H^{3}(\mathbb{A}^{4}_{F}\setminus\{0\},\underline{K}^{\operatorname{MW}}_{2})$ with $\operatorname{W}(F)$ via the (noncanonical) isomorphism $\zeta^{\prime}:H^{3}(\mathbb{A}^{4}_{F}\setminus\{0\},\underline{K}^{\operatorname{MW}}_{2})\to\operatorname{W}(F)$ which is defined in a similar manner to $\zeta$, with $\overline{x}^{*}\wedge\overline{y}^{*}\wedge\overline{z}^{*}\wedge\overline{t}^{*}\in\det(\mathcal{N}_{\{0\}/\mathbb{A}^{4}_{F}})$ instead of $\overline{u}^{*}\wedge\overline{v}^{*}\in\det(\mathcal{N}_{\{0\}/\mathbb{A}^{2}_{F}})$ (see Definition B.19).
@@ 2.3. Invariants of the quadratic linking degree
The rank modulo $2$ of $d_{1}$, the rank modulo $2$ of $d_{2}$ and the rank modulo $2$ of $\operatorname{AQld}_{\mathscr{L}}$ are invariant under changes of orientation classes $\overline{o_{1}},\overline{o_{2}}$ and under changes of parametrisations of $\varphi_{1},\varphi_{2}:\mathbb{A}^{2}_{F}\setminus\{0\}\to X$.
@@
For all $a\in F^{*}$, the rank modulo $2$ of an element of the Witt ring $\operatorname{W}(F)$ is invariant under the multiplication by $\langle a\rangle$. The result follows directly from Lemma 2.16. ∎
@@
In the family of examples 4.3, we provide for each positive integer $n$ an example of an oriented link over $\mathbb{R}$ whose ambient quadratic linking degree has absolute value $n$ (and the same is true of each component of the quadratic linking degree); for examples of oriented links over $\mathbb{R}$ whose (ambient) quadratic linking degree is 00, see Examples 4.2 (with $a<0$ in these examples).
@@
    $\langle 1\rangle+\langle-1\rangle=0$.
@@ 3.1. Assumptions
with $f_{i}$ and $g_{i}$ irreducible. We also assume that the subscheme of $X\setminus Z$ given by the equations $g_{1}=0$ and $g_{2}=0$ is of codimension $2$ in $X\setminus Z$ and that for each generic point $p$ of an irreducible component of this subscheme, $f_{1}$ and $f_{2}$ are units in the residue field $\kappa(p)$.
@@
We define the orientation $o_{(f_{i},g_{i})}$ as the isomorphism $\nu_{Z_{i}}\to\mathcal{O}_{Z_{i}}\otimes\mathcal{O}_{Z_{i}}$ which sends $\overline{f_{i}}^{*}\wedge\overline{g_{i}}^{*}$ to $1\otimes 1$. By Lemma 2.15, there exists $a_{i}\in F^{*}$ such that $\overline{o_{i}}=\overline{o_{(f_{i},g_{i})}\circ(\times a_{i})}=\overline{o_{(a_{i}^{-1}f_{i},g_{i})}}$. Without loss of generality (since we can replace $f_{i}$ with $a_{i}^{-1}f_{i}$), we assume that $o_{i}=o_{(f_{i},g_{i})}$.
@@ 3.3. Computing the quadratic linking class and degree
From Definition B.18 and the notations in Subsection 3.2, using the canonical isomorphism $K^{\operatorname{MW}}_{-2}(F)\simeq\operatorname{W}(F)$ (which sends $\eta^{2}$ to $1$), the final step gives:
@@
From Definition B.19 and the notations in Subsection 3.2, using the canonical isomorphism $K^{\operatorname{MW}}_{-2}(F)\simeq\operatorname{W}(F)$ (which sends $\eta^{2}$ to $1$), the ambient quadratic linking degree of $\mathscr{L}$ is equal to:
@@ 4. Examples of computations of the quadratic linking degree
  • $Z_{1}$ is the intersection of the closed subscheme of $\mathbb{A}^{4}_{F}=\operatorname{Spec}(F[x,y,z,t])$ of ideal $(x,y)$ and of $X:=\mathbb{A}^{4}_{F}\setminus\{0\}$;
  • $\varphi_{1}:\mathbb{A}^{2}_{F}\setminus\{0\}\to X$ is the morphism associated to the morphism of $F$-algebras $F[x,y,z,t]\to F[u,v]$ which maps $x,y,z,t$ to $0,0,u,v$ respectively;
  • $\overline{o_{1}}$ is the orientation class associated to the couple $(x,y)$ (i.e. the class of the isomorphism $o_{1}:\nu_{Z_{1}}\to\mathcal{O}_{Z_{1}}\otimes\mathcal{O}_{Z_{1}}$ which maps $\overline{x}^{*}\wedge\overline{y}^{*}$ to $1\otimes 1$);
  • $Z_{2}$ is the intersection of the closed subscheme of $\mathbb{A}^{4}_{F}$ of ideal $(z,t)$ and of $X$;
@@
  • $\overline{o_{2}}$ is the orientation class associated to the couple $(z,t)$ (i.e. the class of the isomorphism $o_{2}:\nu_{Z_{2}}\to\mathcal{O}_{Z_{2}}\otimes\mathcal{O}_{Z_{2}}$ which maps $\overline{z}^{*}\wedge\overline{t}^{*}$ to $1\otimes 1$).
@@
Note that the rank modulo $2$ of each component of the quadratic linking degree and of the ambient quadratic linking degree of the Hopf link is equal to $1$. Note that for every positive even integer $k$, the image by $\Sigma_{k}$ of each component of the quadratic linking degree and of the ambient quadratic linking degree of the Hopf link is equal to 00. Note that if $F=\mathbb{R}$ then the absolute value of each component of the quadratic linking degree and of the ambient quadratic linking degree of the Hopf link is equal to $1$.
@@
Quadratic linking class $-\langle z\rangle\eta\otimes(\overline{t}^{*}\wedge\overline{x}^{*}\wedge\overline{y}^{*})$ $\oplus$ $\langle x\rangle\eta\otimes(\overline{y}^{*}\wedge\overline{z}^{*}\wedge\overline{t}^{*})$
Apply $\widetilde{o_{1}}\oplus\widetilde{o_{2}}$ $-\langle z\rangle\eta\otimes\overline{t}^{*}$ $\oplus$ $\langle x\rangle\eta\otimes\overline{y}^{*}$
Apply $\varphi_{1}^{*}\oplus\varphi_{2}^{*}$ $-\langle u\rangle\eta\otimes\overline{v}^{*}$ $\oplus$ $\langle u\rangle\eta\otimes\overline{v}^{*}$
@@
Quadratic linking degree $-1$ $\oplus$ $1$
Quadratic linking class $-\langle z\rangle\eta\otimes(\overline{t}^{*}\wedge\overline{x}^{*}\wedge\overline{y}^{*})$ $\oplus$ $\langle x\rangle\eta\otimes(\overline{y}^{*}\wedge\overline{z}^{*}\wedge\overline{t}^{*})$
@@
Ambient quadratic linking degree $-1$
@@
(Binary links) Let $F$ be a perfect field of characteristic different from $2$ and $a\in F^{*}$. We define the binary link $B_{a}$ over $F$ as follows:
@@
Note that the rank modulo $2$ of each component of the quadratic linking degree and of the ambient quadratic linking degree of the binary link $B_{a}$ is 00 (hence the invariant presented in Corollary 2.17 distinguishes between the Hopf link and the binary links). Note that the image by $\Sigma_{2}$ of each component of the quadratic linking degree and of the ambient quadratic linking degree of the binary link $B_{a}$ is $\langle a\rangle\in\operatorname{W}(F)/(1)$. For instance, if $F=\mathbb{Q}$, $\Sigma_{2}$ distinguishes between all the $B_{p}$ with $p$ prime numbers since if $p\neq q$ are prime numbers then $\langle p\rangle\in\operatorname{W}(\mathbb{Q})/(1)$ corresponds to $1\in\operatorname{W}(\mathbb{Z}/p\mathbb{Z})\subset\bigoplus_{r\text{ prime}}\operatorname{W}(\mathbb{Z}/r\mathbb{Z})$ and $\langle q\rangle\in\operatorname{W}(\mathbb{Q})/(1)$ corresponds to $1\in\operatorname{W}(\mathbb{Z}/q\mathbb{Z})\subset\bigoplus_{r\text{ prime}}\operatorname{W}(\mathbb{Z}/r\mathbb{Z})$ (via the isomorphism $\operatorname{W}(\mathbb{Q})/(1)\simeq\bigoplus_{r\text{ prime}}\operatorname{W}(\mathbb{Z}/r\mathbb{Z})$ induced by the isomorphism $\operatorname{W}(\mathbb{Q})\simeq\operatorname{W}(\mathbb{R})\oplus\bigoplus_{r\text{ prime}}\operatorname{W}(\mathbb{Z}/r\mathbb{Z})$ which maps $\langle p_{1}\dots p_{n}\rangle\in\operatorname{W}(\mathbb{Q})$ (with $p_{1},\dots,p_{n}$ distinct primes) to $\langle p_{1}\dots p_{n}\rangle\oplus\bigoplus_{i=1}^{n}\langle\prod_{j\neq i}p_{j}\rangle\in\operatorname{W}(\mathbb{R})\oplus\bigoplus_{i=1}^{n}\operatorname{W}(\mathbb{Z}/p_{i}\mathbb{Z})$). Note that if $F=\mathbb{R}$ then the absolute value of each component of the quadratic linking degree and of the ambient quadratic linking degree of the binary link $B_{a}$ is equal to $2$ if $a>0$, to 00 if $a<0$ (hence the invariant presented in Corollary 2.18 distinguishes between the Hopf link and the binary links, as well as between the binary links with positive parameter and the binary links with negative parameter).
@@
Quad. link. class $-\langle f_{2}\rangle\eta\otimes(\overline{g_{2}}^{*}\wedge\overline{f_{1}}^{*}\wedge\overline{g_{1}}^{*})\cdot(x-y)$ $\oplus$ $\langle f_{1}\rangle\eta\otimes(\overline{g_{1}}^{*}\wedge\overline{f_{2}}^{*}\wedge\overline{g_{2}}^{*})\cdot(x-y)$
@@
Apply $\widetilde{o_{1}}\oplus\widetilde{o_{2}}$ $-\langle f_{2}\rangle\eta\otimes\overline{g_{2}}^{*}\cdot(x-y)$ $\oplus$ $\phantom{+}\langle f_{1}\rangle\eta\otimes\overline{g_{1}}^{*}\cdot(x-y)$
@@
Apply $\varphi_{1}^{*}\oplus\varphi_{2}^{*}$ $-\langle au\rangle\eta\otimes\overline{u-v}^{*}\cdot(u-v)$ $\oplus$ $\phantom{+}\langle au\rangle\eta\otimes\overline{u-v}^{*}\cdot(u-v)$
@@
Quad. lk. deg. $1+\langle a\rangle$ $\oplus$ $-(1+\langle a\rangle)$
Quad. link. class $-\langle f_{2}\rangle\eta\otimes(\overline{g_{2}}^{*}\wedge\overline{f_{1}}^{*}\wedge\overline{g_{1}}^{*})\cdot(x-y)$ $\oplus$ $\langle f_{1}\rangle\eta\otimes(\overline{g_{1}}^{*}\wedge\overline{f_{2}}^{*}\wedge\overline{g_{2}}^{*})\cdot(x-y)$
@@
The following family of examples is an analogue of the family of torus links $T(2,2n)$ (with $n\geq 1$ an integer) in knot theory. Note that $T(2,2)$ is the Hopf link and that its analogue below is slightly different from the Hopf link in the example above and has quadratic linking degree $(1,-1)$ and ambient quadratic linking degree $-1$. Note that $T(2,4)$ is the Solomon link.
@@
Recall that (in knot theory) one of the components of $T(2,2n)$ is the intersection of $\{(a,b)\in\mathbb{C}^{2},b=a^{n}\}$ with $\mathbb{S}^{3}_{\varepsilon}$, the $3$-sphere of radius $\varepsilon$, and that the other component of $T(2,2n)$ is the intersection of $\{(a,b)\in\mathbb{C}^{2},b=-a^{n}\}$ with $\mathbb{S}^{3}_{\varepsilon}$ (for $\varepsilon>0$ small enough). By writing $a=x+iy$ and $b=z+it$ (with $x,y,z,t\in\mathbb{R}$), the equation $b=a^{n}$ becomes the system of equations
@@
Note that the rank modulo $2$ of each component of the quadratic linking degree and of the ambient quadratic linking degree of the analogue of $T(2,2n)$ is $1$ if $n$ is odd, 00 if $n$ is even. Note that the absolute value of each component of the quadratic linking degree and of the ambient quadratic linking degree of the analogue of $T(2,2n)$ is equal to $n$, hence the invariant presented in Corollary 2.18 distinguishes between all these links $T(2,2n)$, similarly to the absolute value of the linking number which distinguishes between all the links $T(2,2n)$ in knot theory (recall that the linking number of $T(2,2n)$ is equal to $n$ in classical knot theory).

### v6

@@ 1. Introduction
We begin by defining counterparts, over a perfect field $F$, to oriented links with two components (i.e. couples of disjoint oriented knots). Specifically, we replace the circle $\mathbb{S}^{1}$ with $\mathbb{A}^{2}_{F}\setminus\{0\}$ and the 333-sphere $\mathbb{S}^{3}$ with $\mathbb{A}^{4}_{F}\setminus\{0\}$. We then define a counterpart to the linking number, which in knot theory is an invariant of oriented links with two components which corresponds to the number of times one of the oriented components turns around the other oriented component, and a counterpart to the linking couple, which is a couple of integers whose absolute values coincide with the absolute value of the linking number. We call these counterparts the ambient quadratic linking degree and the quadratic linking degree (couple) respectively. To define these, we use Chow-Witt groups, and more generally Rost-Schmid groups, instead of the singular cohomology groups used in classical knot theory. The Rost-Schmid groups are the cohomology groups of the Rost-Schmid complex (see Appendix B) which is built from Milnor-Witt $K$-theory groups; this is where motivic homotopy theory comes into play, as Milnor-Witt $K$-theory groups are in fact groups of morphisms of motivic spheres in the stable motivic homotopy category (see [Mor12, Corollary 1.25]).
@@
In Section 2, we give the definitions of the quadratic linking class, the quadratic linking degree and the ambient quadratic linking degree, then we determine how they depend on choices of orientations and of parametrisations of $\mathbb{A}^{2}_{F}\setminus\{0\}\to\mathbb{A}^{4}_{F}\setminus\{0\}$ (see Lemma 2.16) and deduce invariants of the (ambient) quadratic linking degree (see Corollaries 2.17 and 2.18 and Theorem 2.21). For instance, in the case $F=\mathbb{R}$, the absolute values of the components of the quadratic linking degree and of the ambient quadratic linking degree (which are in $\operatorname{W}(\mathbb{R})\simeq\mathbb{Z}$) are invariant under changes of orientations and of parametrisations of $\mathbb{A}^{2}_{F}\setminus\{0\}\to\mathbb{A}^{4}_{F}\setminus\{0\}$. This is similar to the fact that the absolute value of the linking number does not depend on choices of orientations. In the general case, the ranks modulo 222 of the components of the quadratic linking degree and of the ambient quadratic linking degree are invariants, and more importantly we have the following lemma-definition and theorem:
@@
In Section 3 we give a method to explicitly compute the quadratic linking class (see Theorem 3.1), the quadratic linking degree (see Theorem 3.2) and the ambient quadratic linking degree (see Theorem 3.3) when the link $Z_{1}\sqcup Z_{2}\subset\mathbb{A}^{4}_{F}\setminus\{0\}$ is such that for each $i\in\{1,2\}$ the closure $\overline{Z_{i}}\subset\mathbb{A}^{4}_{F}$ of $Z_{i}$ is given by two irreducible equations $\{f_{i}=0,g_{i}=0\}$ such that $\{g_{1}=0,g_{2}=0\}$ is of codimension 222 in $(\mathbb{A}^{4}_{F}\setminus\{0\})\setminus(Z_{1}\sqcup Z_{2})$ and for each generic point $p$ of an irreducible component of $\{g_{1}=0,g_{2}=0\}$, $f_{1}$ and $f_{2}$ are units in the residue field $\kappa(p)$.
@@ 2.1. Conventions and notations
where $\partial$ is the boundary map (see Definition B.12 and Theorem B.13) and the map in the middle is induced by the isomorphism $\det(\mathcal{N}_{\{0\}/\mathbb{A}^{2}_{F}})\to\mathcal{O}_{\{0\}}\otimes\mathcal{O}_{\{0\}}$ which sends $\overline{u}^{*}\wedge\overline{v}^{*}$ to $\otimes 1$ 1 (see Definition B.18). We also identify $H^{3}(\mathbb{A}^{4}_{F}\setminus\{0\},\underline{K}^{\operatorname{MW}}_{2})$ with $\operatorname{W}(F)$ via the (noncanonical) isomorphism $\zeta^{\prime}:H^{3}(\mathbb{A}^{4}_{F}\setminus\{0\},\underline{K}^{\operatorname{MW}}_{2})\to\operatorname{W}(F)$ which is defined in a similar manner to $\zeta$, with $\overline{x}^{*}\wedge\overline{y}^{*}\wedge\overline{z}^{*}\wedge\overline{t}^{*}\in\det(\mathcal{N}_{\{0\}/\mathbb{A}^{4}_{F}})$ instead of $\overline{u}^{*}\wedge\overline{v}^{*}\in\det(\mathcal{N}_{\{0\}/\mathbb{A}^{2}_{F}})$ (see Definition B.19).
@@ 2.3. Invariants of the quadratic linking degree
The rank modulo 222 of $d_{1}$, the rank modulo 222 of $d_{2}$ and the rank modulo 222 of $\operatorname{AQld}_{\mathscr{L}}$ are invariant under changes of orientation classes $\overline{o_{1}},\overline{o_{2}}$ and under changes of parametrisations of $\varphi_{1},\varphi_{2}:\mathbb{A}^{2}_{F}\setminus\{0\}\to X$.
@@
For all $a\in F^{*}$, the rank modulo 222 of an element of the Witt ring $\operatorname{W}(F)$ is invariant under the multiplication by $\langle a\rangle$. The result follows directly from Lemma 2.16. ∎
@@
In the family of examples 4.3, we provide for each positive integer $n$ an example of an oriented link over $\mathbb{R}$ whose ambient quadratic linking degree has absolute value $n$ (and the same is true of each component of the quadratic linking degree); for examples of oriented links over $\mathbb{R}$ whose (ambient) quadratic linking degree is 0, see Examples 4.2 (with $a<0$ in these examples).
@@
    1+1=0$\langle$ 1$\rangle$+$\langle$-1$\rangle$=0⟨ 1 ⟩ + ⟨ - 1 ⟩ = 0.
@@ 3.1. Assumptions
with $f_{i}$ and $g_{i}$ irreducible. We also assume that the subscheme of $X\setminus Z$ given by the equations $g_{1}=0$ and $g_{2}=0$ is of codimension 222 in $X\setminus Z$ and that for each generic point $p$ of an irreducible component of this subscheme, $f_{1}$ and $f_{2}$ are units in the residue field $\kappa(p)$.
@@
We define the orientation $o_{(f_{i},g_{i})}$ as the isomorphism $\nu_{Z_{i}}\to\mathcal{O}_{Z_{i}}\otimes\mathcal{O}_{Z_{i}}$ which sends $\overline{f_{i}}^{*}\wedge\overline{g_{i}}^{*}$ to $\otimes 1$ 1. By Lemma 2.15, there exists $a_{i}\in F^{*}$ such that $\overline{o_{i}}=\overline{o_{(f_{i},g_{i})}\circ(\times a_{i})}=\overline{o_{(a_{i}^{-1}f_{i},g_{i})}}$. Without loss of generality (since we can replace $f_{i}$ with $a_{i}^{-1}f_{i}$), we assume that $o_{i}=o_{(f_{i},g_{i})}$.
@@ 3.3. Computing the quadratic linking class and degree
From Definition B.18 and the notations in Subsection 3.2, using the canonical isomorphism $K^{\operatorname{MW}}_{-2}(F)\simeq\operatorname{W}(F)$ (which sends $\eta^{2}$ to 111), the final step gives:
@@
From Definition B.19 and the notations in Subsection 3.2, using the canonical isomorphism $K^{\operatorname{MW}}_{-2}(F)\simeq\operatorname{W}(F)$ (which sends $\eta^{2}$ to 111), the ambient quadratic linking degree of $\mathscr{L}$ is equal to:
@@ 4. Examples of computations of the quadratic linking degree
[…]
@@
[…]
@@
Note that the rank modulo 222 of each component of the quadratic linking degree and of the ambient quadratic linking degree of the Hopf link is equal to 111. Note that for every positive even integer $k$, the image by $\Sigma_{k}$ of each component of the quadratic linking degree and of the ambient quadratic linking degree of the Hopf link is equal to 0. Note that if $F=\mathbb{R}$ then the absolute value of each component of the quadratic linking degree and of the ambient quadratic linking degree of the Hopf link is equal to 111.
@@
Quadratic linking class $-\langle z\rangle\eta\otimes(\overline{t}^{*}\wedge\overline{x}^{*}\wedge\overline{y}^{*})$oplus $\langle x\rangle\eta\otimes(\overline{y}^{*}\wedge\overline{z}^{*}\wedge\overline{t}^{*})$
Apply $\widetilde{o_{1}}\oplus\widetilde{o_{2}}$ $-\langle z\rangle\eta\otimes\overline{t}^{*}$oplus $\langle x\rangle\eta\otimes\overline{y}^{*}$
Apply $\varphi_{1}^{*}\oplus\varphi_{2}^{*}$ $-\langle u\rangle\eta\otimes\overline{v}^{*}$oplus $\langle u\rangle\eta\otimes\overline{v}^{*}$
@@
Quadratic linking degree 1-1- 1 $\oplus$ 111
Quadratic linking class $-\langle z\rangle\eta\otimes(\overline{t}^{*}\wedge\overline{x}^{*}\wedge\overline{y}^{*})$oplus $\langle x\rangle\eta\otimes(\overline{y}^{*}\wedge\overline{z}^{*}\wedge\overline{t}^{*})$
@@
Ambient quadratic linking degree 1-1- 1
@@
(Binary links) Let $F$ be a perfect field of characteristic different from 222 and $a\in F^{*}$. We define the binary link $B_{a}$ over $F$ as follows:
@@
Note that the rank modulo 222 of each component of the quadratic linking degree and of the ambient quadratic linking degree of the binary link $B_{a}$ is 0 (hence the invariant presented in Corollary 2.17 distinguishes between the Hopf link and the binary links). Note that the image by $\Sigma_{2}$ of each component of the quadratic linking degree and of the ambient quadratic linking degree of the binary link $B_{a}$ is $\langle a\rangle\in\operatorname{W}(F)/(1)$. For instance, if $F=\mathbb{Q}$, $\Sigma_{2}$ distinguishes between all the $B_{p}$ with $p$ prime numbers since if $p\neq q$ are prime numbers then $\langle p\rangle\in\operatorname{W}(\mathbb{Q})/(1)$ corresponds to $1\in\operatorname{W}(\mathbb{Z}/p\mathbb{Z})\subset\bigoplus_{r\text{ prime}}\operatorname{W}(\mathbb{Z}/r\mathbb{Z})$ and $\langle q\rangle\in\operatorname{W}(\mathbb{Q})/(1)$ corresponds to $1\in\operatorname{W}(\mathbb{Z}/q\mathbb{Z})\subset\bigoplus_{r\text{ prime}}\operatorname{W}(\mathbb{Z}/r\mathbb{Z})$ (via the isomorphism $\operatorname{W}(\mathbb{Q})/(1)\simeq\bigoplus_{r\text{ prime}}\operatorname{W}(\mathbb{Z}/r\mathbb{Z})$ induced by the isomorphism $\operatorname{W}(\mathbb{Q})\simeq\operatorname{W}(\mathbb{R})\oplus\bigoplus_{r\text{ prime}}\operatorname{W}(\mathbb{Z}/r\mathbb{Z})$ which maps $\langle p_{1}\dots p_{n}\rangle\in\operatorname{W}(\mathbb{Q})$ (with $p_{1},\dots,p_{n}$ distinct primes) to $\langle p_{1}\dots p_{n}\rangle\oplus\bigoplus_{i=1}^{n}\langle\prod_{j\neq i}p_{j}\rangle\in\operatorname{W}(\mathbb{R})\oplus\bigoplus_{i=1}^{n}\operatorname{W}(\mathbb{Z}/p_{i}\mathbb{Z})$). Note that if $F=\mathbb{R}$ then the absolute value of each component of the quadratic linking degree and of the ambient quadratic linking degree of the binary link $B_{a}$ is equal to 222 if $a>0$, to 0 if $a<0$ (hence the invariant presented in Corollary 2.18 distinguishes between the Hopf link and the binary links, as well as between the binary links with positive parameter and the binary links with negative parameter).
@@
Quad. link. class $-\langle f_{2}\rangle\eta\otimes(\overline{g_{2}}^{*}\wedge\overline{f_{1}}^{*}\wedge\overline{g_{1}}^{*})\cdot(x-y)$oplus $\langle f_{1}\rangle\eta\otimes(\overline{g_{1}}^{*}\wedge\overline{f_{2}}^{*}\wedge\overline{g_{2}}^{*})\cdot(x-y)$
@@
Apply $\widetilde{o_{1}}\oplus\widetilde{o_{2}}$ $-\langle f_{2}\rangle\eta\otimes\overline{g_{2}}^{*}\cdot(x-y)$oplus $\phantom{+}\langle f_{1}\rangle\eta\otimes\overline{g_{1}}^{*}\cdot(x-y)$
@@
Apply $\varphi_{1}^{*}\oplus\varphi_{2}^{*}$ $-\langle au\rangle\eta\otimes\overline{u-v}^{*}\cdot(u-v)$oplus $\phantom{+}\langle au\rangle\eta\otimes\overline{u-v}^{*}\cdot(u-v)$
@@
Quad. lk. deg. $1+\langle a\rangle$oplus $-(1+\langle a\rangle)$
Quad. link. class $-\langle f_{2}\rangle\eta\otimes(\overline{g_{2}}^{*}\wedge\overline{f_{1}}^{*}\wedge\overline{g_{1}}^{*})\cdot(x-y)$oplus $\langle f_{1}\rangle\eta\otimes(\overline{g_{1}}^{*}\wedge\overline{f_{2}}^{*}\wedge\overline{g_{2}}^{*})\cdot(x-y)$
@@
The following family of examples is an analogue of the family of torus links $T(2,2n)$ (with $n\geq 1$ an integer) in knot theory. Note that $T(2,2)$ is the Hopf link and that its analogue below is slightly different from the Hopf link in the example above and has quadratic linking degree (1,1)(1,-1)( 1 , - 1 ) and ambient quadratic linking degree 1-1- 1. Note that $T(2,4)$ is the Solomon link.
@@
Recall that (in knot theory) one of the components of $T(2,2n)$ is the intersection of $\{(a,b)\in\mathbb{C}^{2},b=a^{n}\}$ with $\mathbb{S}^{3}_{\varepsilon}$, the 333-sphere of radius $\varepsilon$, and that the other component of $T(2,2n)$ is the intersection of $\{(a,b)\in\mathbb{C}^{2},b=-a^{n}\}$ with $\mathbb{S}^{3}_{\varepsilon}$ (for $\varepsilon>0$ small enough). By writing $a=x+iy$ and $b=z+it$ (with $x,y,z,t\in\mathbb{R}$), the equation $b=a^{n}$ becomes the system of equations
@@
Note that the rank modulo 222 of each component of the quadratic linking degree and of the ambient quadratic linking degree of the analogue of $T(2,2n)$ is 111 if $n$ is odd, 0 if $n$ is even. Note that the absolute value of each component of the quadratic linking degree and of the ambient quadratic linking degree of the analogue of $T(2,2n)$ is equal to $n$, hence the invariant presented in Corollary 2.18 distinguishes between all these links $T(2,2n)$, similarly to the absolute value of the linking number which distinguishes between all the links $T(2,2n)$ in knot theory (recall that the linking number of $T(2,2n)$ is equal to $n$ in classical knot theory).